# Effects of model misspecification on small area estimators

### v1

@@ 3 Conclusion
In this paper, we investigate the effects of misspecified mean structure and sampling variance in the well-known nested error regression model on EBLUP and OBP that uses (i) unit-level auxiliary variables only and (ii) area-level auxiliary variables only (i.e., unit context model). Through simulated scenarios, we illustrate that, in the presence of substantial model misspecifications, OBP procedure for unit context model shows better performance than that of the corresponding unit-level model using unit-specific auxiliary variables. Also, with known sampling variances, the performance of OBP under area-level model and unit-context model is similar, and both are better than the OBP using unit-specific auxiliary variables. These findings suggest that implementing the OBP procedure within unit-context models could be a viable alternative when faced with challenges such as the absence of census information and/or potential model misspecification issues. While we do not have a clear understanding why OBP for nested error model with solely unit level auxiliary variables performs poorly in our simulation study, it appears that the difference between the sample and population means of the auxiliary variables could be a potential factor. We encourage further theoretical and empirical research to understand this unexpected behavior of OBP for nested error model with solely unit level auxiliary variables.

### v2

@@ 3 Simulations
Remark 2: Minimizing $Q(\bm{\psi})$ vs Minimizing $\rm{MSPE}(\tilde{\theta}(\bm{\psi}))$
Another possible explanation for the underperformance of OBP-UNIT might be that the basic OBP equation, valid for fixed $\bm{\bm{\psi}}$, may not hold for random $\bm{\bm{\psi}}$, as in the case of BPE or MLE. To investigate this, we begin with the basic OBP equation:
$\rm{E}(|\tilde{\bm{\theta}}(\bm{\psi})-\bm{\theta}|^{2})=\rm{E}\{Q(\bm{\psi})\}.$ (3.9)
Let $\hat{\bm{\psi}}$ be the BPE of $\bm{\psi}$ obtained by minimizing the observed MSPE, and let $\tilde{\bm{\psi}}$ denote the MLE. The corresponding OBP-UNIT and EBLUP-UNIT are $\tilde{\bm{\theta}}(\hat{\bm{\psi}})$ and $\tilde{\bm{\theta}}(\tilde{\bm{\psi}})$, respectively. If the equation (3.9) holds for random $\bm{\psi}$, one would have
$\rm{E}(|\tilde{\bm{\theta}}(\hat{\bm{\psi}})-\bm{\theta}|^{2})=\rm{E}\{Q(\hat{\bm{\psi}})\}\leq\rm{E}\{Q(\tilde{\bm{\psi}})\}=\rm{E}(|\tilde{\bm{\theta}}(\tilde{\bm{\psi}})-\bm{\theta}|^{2}).$ (3.10)
The inequality follows from the definition of the BPE. To test this, we compare the four terms in the inequality using the same simulation settings as in previous analyses. The results are presented in Tables 3.3 and 3.4.
When the assumed model is completely different from the true model, Table 3.3 indicates that $\rm{E}(|\tilde{\bm{\theta}}(\hat{\bm{\psi})}-\bm{\theta}|^{2})$, the first term of (3.10), could be much larger than $\rm{E}\{Q(\hat{\bm{\psi}})\}$, the second term of (3.10), unless $m$ is extremely large. But $\rm{E}\{Q(\tilde{\bm{\psi}})\}$, third term of (3.10) is approximately equal to $\rm{E}(|\tilde{\bm{\theta}}(\tilde{\bm{\psi}})-\bm{\theta}|^{2})$, the third term of (3.10) even for smaller $m$. When the assumed model is partially correct, Table 3.4 shows neither $\rm{E}(|\tilde{\bm{\theta}}(\hat{\bm{\psi}})-\bm{\theta}|^{2})=\rm{E}\{Q(\hat{\bm{\psi}})\}$ nor $\rm{E}\{Q(\tilde{\bm{\psi}})\}=\rm{E}(|\tilde{\bm{\theta}}(\tilde{\bm{\psi}})-\bm{\theta}|^{2})$ of (3.10) holds, with $\rm{E}(|\tilde{\bm{\theta}}(\hat{\bm{\psi}})-\bm{\theta}|^{2})$ being much larger than $\rm{E}{Q(\hat{\bm{\psi}})}$, especially for smaller $m$ (e.g., 40). Thus our simulation supports the argument that MSPE of OBP-UNIT could be larger than that of MSPE of EBLUP-UNIT unless $m$ is extremely and unreasonably large.
One potential factor contributing to these differences between $\rm{E}(|\tilde{\bm{\theta}}(\hat{\bm{\psi}})-\bm{\theta}|^{2})$ and $\rm{E}\{Q(\hat{\bm{\psi}})\}$ is the numerical stability of $\hat{\bm{\psi}}$, BPE of $\bm{\psi}$. Particularly when $m=40$, nearly 20%, of BPE estimates yielding a zero value for the ratio $\sigma^{2}_{u}/\sigma^{2}_{e}$. The performance of OBP-UNIT appears to depend on value of $m$. The performance of OBP-UNIT improves for extremely large $m$ (e.g., 4000 or 40000), where both sides of the inequality of (3.10), i.e., $\rm{E}(|\tilde{\bm{\theta}}(\hat{\bm{\psi}})-\bm{\theta}|^{2})=\rm{E}\{Q(\hat{\bm{\psi}})\}$ and $\rm{E}\{Q(\tilde{\bm{\psi}})\}=\rm{E}(|\tilde{\bm{\theta}}(\tilde{\bm{\psi}})-\bm{\theta}|^{2})$, tend to be approximately hold, and thus OBP-UNIT eventually perform marginally better than EBLUP-UNIT under the model misspecifications. We hypothesize that the BPE of variance components stabilizes as $m$ increases, suggesting that future research should work on improving the BPE of variance components before proceeding the OBP method under unit-level models.
Table 3.3: Comparison among the four terms in (3.10) under Table 3.1 settings
$(m,b)$ $\rm{MSPE}(\tilde{\bm{\theta}}(\hat{\bm{\psi}}))$ $\rm{E}\{Q(\hat{\bm{\psi}})\}$ $\rm{E}\{Q(\tilde{\bm{\psi}})\}$ $\rm{MSPE}(\tilde{\bm{\theta}}(\tilde{\bm{\psi}}))$
(40, 5) 1.775 0.448 1.479 1.474
(100, 5) 1.409 0.918 1.493 1.495
(400, 5) 1.352 1.211 1.500 1.508
(1000, 5) 1.384 1.323 1.504 1.513
(2000, 5) 1.414 1.379 1.521 1.522
(4000, 5) 1.414 1.398 1.512 1.515
(40, 10) 3.919 -1.189 1.649 1.522
(100, 10) 2.364 0.493 1.551 1.513
(400, 10) 1.711 1.231 1.508 1.510
(1000, 10) 1.600 1.389 1.530 1.532
(2000, 10) 1.509 1.412 1.500 1.496
(4000, 10) 1.485 1.431 1.493 1.493
Table 3.4: Comparison among the four terms in (3.10) under Table 3.2 settings
$m$ $\rm{MSPE}(\tilde{\bm{\theta}}(\hat{\bm{\psi}}))$ $\rm{E}\{Q(\hat{\bm{\psi}})\}$ $\rm{E}\{Q(\tilde{\bm{\psi}})\}$ $\rm{MSPE}(\tilde{\bm{\theta}}(\tilde{\bm{\psi}}))$
40 18.230 -14.036 3.247 1.532
100 7.958 -4.809 1.586 1.514
400 3.122 -0.584 1.583 1.508
1000 2.173 0.828 1.580 1.569
2000 1.815 1.131 1.546 1.501
4000 1.650 1.267 1.496 1.505
40000 1.488 1.408 1.462 1.503
@@ 4 Conclusion
In this paper, we investigate the effects of misspecified mean structure and sampling variance in the well-known nested error regression model on EBLUP and OBP that uses (i) unit-level auxiliary variables only and (ii) area-level auxiliary variables only (i.e., unit context model). Through a series of simulations, we demonstrate that under significant model misspecifications, the OBP procedure for the area-level model and the unit context model outperforms the corresponding unit-level model that relies solely on unit-specific auxiliary variables.
We propose two potential reasons for the underperformance of OBP-UNIT: first, the difference between the sample and population means of the auxiliary variables may negatively affect the performance of OBP-UNIT; second, the numerical stability of the BPE for variance components using unit-level auxiliary variables could be a contributing factor. Our results suggest that utilizing the OBP procedure with area-level auxiliary variables is a promising alternative, particularly when challenges such as the lack of census information and model misspecification is a concern.
Furthermore, when choosing between the unit-context model and the area-level model, we recommend the area-level model as it may offer greater stability. In the area-level model, the variance of sampling errors can be estimated using smoothing methods and is less dependent on the assumed model which may be misspecified. Finally, we recommend further theoretical and empirical research to explore and better understand these unexpected outcomes in the OBP for nested error models that utilize only unit-level auxiliary variables.